OpenStudy (anonymous):

Solve the following system. y = (1/2)x^2 + 2x - 1 and 3x - y = 1 The solutions are ( ) and ( ) (remember to include the commas)

OpenStudy (danjs):

i would solve the second one for y = , then just set them equal y = y

OpenStudy (danjs):

solve for the x value(s) and then use those x's to get the corresponding y value(s)

OpenStudy (danjs):

y = (1/2)x^2 + 2x - 1 and y = 3x - 1 solve, 3x - 1 = (1/2)x^2 + 2x - 1

OpenStudy (anonymous):

which when you combine like terms will give you 1/2x^2-x = 0

OpenStudy (anonymous):

I got x=2

OpenStudy (anonymous):

then you can just factor it using the quadratic formula. Yes you should get X = 2 and X = 0